(x^2)+25=10x
What steps do I need to do to find x? It has been explained to me that I need to "factor" something out, but I am not sure what or any kinds of steps I need to take to even start solving. Thanks.

Answer by ankor@dixie-net.com(22740)

(x^2)+25=10x
arrange as a quadratic equation, subtract 10x from both sides and you have:
x^2 - 10x + 25 = 0
Think what are the factors of 25, that add up to -10,
that would be -5 * -5, right? So our factors are:
(x-5)(x-5) = 0; Prove it to yourself, FOIL (x-5)(x-5)
therefore
x = +5